A transformer rated at a maximum power of 10 kW is used to connect a 5000 V transmission line to a 240 V circuit. (a) Find the ratio of turns in the windings of the transformer. [4] (b) Calculate the maximum current in the 240 V circuit.
1
Expert's answer
2019-08-19T09:49:26-0400

(a) The ratio of turns in the windings of the transformer


"\frac{N_1}{N_2}=\frac{V_1}{V_2}=\frac{5000}{240}=\frac{125}{6}=20.83"

(b)


"P_1=P_2=10000\:\rm{W}"

So, the maximum current in the secondary coil


"I_2=\frac{P_2}{V_2}=\frac{10000\:\rm{W}}{240\:\rm{V}}=41.67\:\rm{A}"